two electrostatic point charges of +60.0 μc and +50.0 μc exert a repulsive force on each other of 175 n. what is the distance between the two charges?


1
Expert's answer
2021-08-11T09:53:29-0400

The Coulomb's law says

"F=k\frac{q_1q_2}{r^2}"

Hence, the distance between two charges

"r=\sqrt{\frac{kq_1q_2}{F}}\\=\sqrt{\frac{9*10^9*60*10^{-6}*50*10^{-6}}{175}}=0.393\:\rm m"